# Photonic Exceptional Points in Holography and QCD

### v4

@@ 3.1 The holographic lasing spectra of various pumping patterns
Now, based on our model of three coupled flavor-brane gauge modes $a_{i}(z)$, we can extend the work of [41] and find analytical results for several symmetric cases: gain-gain-gain, gain-loss-loss, and gain-gain-loss patterns, with partially pumped or evenly pumped structures, PT-symmetric gain-loss arrangements (unbroken to broken), and identify EP2/EP3.
@@
Figure 4: Two runs for finding eigenvalue trajectories in the complex plane as $\gamma$ is varied for the ternary coupled microcavity model (anti-PT sweep). In the left $N_{z}=20$ (grid points along holographic z) and in the right $N_{z}=200$.
Figure 5: Two runs for finding imaginary parts versus $\gamma$ (growth rates of modes) for the ternary coupled microcavity model (anti-PT sweep). In the left $N_{z}=20$ (grid points along holographic z) and in the right $N_{z}=200$.
Figure 6: Two runs for finding radial profiles of the dominant lasing eigenmode at the $\gamma$ where the growth is maximal for the ternary coupled microcavity model (anti-PT sweep). In the left $N_{z}=20$ (grid points along holographic z) and in the right $N_{z}=200$. One could see that similar to confining models, most of the modes are accumulated close to the wall, around $z=0$.
@@
For the even pumping (gain-gain-gain), we have $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma$. For the partial pumping/gain-loss-loss, we get $\gamma_{1}=\gamma$, $\gamma_{2}=-\gamma$, $\gamma_{3}=-\gamma$, or $\gamma_{1}=\gamma$, $\gamma_{2}=\gamma$, $\gamma_{3}=0$. For the anti-PT ternary system, we get $\gamma_{1}=\gamma$, $\gamma_{2}=0$ and $\gamma_{3}=-\gamma$. For the PT pair plus a spectator case, we get a PT pair on modes 1 and 2, while the mode 3 is neutral or weakly coupled. So we have $\gamma_{1}=\gamma$, $\gamma_{2}=-\gamma$, $\gamma_{3}=0$.
@@
Figure 7: Phase diagram distinguishing the unbroken PT phase, where all eigenvalues are real, from the broken PT phase, where at least one eigenvalue is complex.
In figure 7, a phase diagram heatmap is shown. In the blue region (0), there is an unbroken PT phase where all eigenvalues are real, corresponding to symmetric lasing. In the red region (1), the PT symmetry is broken, with complex-conjugate eigenvalue pairs indicating symmetry breaking. This illustrates where the exceptional points occur as transitions between the phases are consistent with the physics of ternary microrings in [42].
@@ 3.2 Analytic eigenvalues for symmetric coupling
For even pumping corresponding to gain–gain–gain, where $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma$, and for the nondegenerate symmetric mode, we get $(1,1,1):\lambda_{+}=i\gamma+2\kappa$, while for the doubly degenerate orthogonal subspace with degeneracy 2, we have $\lambda_{-}=i\gamma-\kappa$.
@@
For the fully symmetric anti-PT diagonal pattern $(i\gamma,0,-i\gamma)$ and uniform $\kappa$, the characteristic cubic must be solved. In this case, the eigenvalues are generally distinct complex numbers, and there is no triple root for nonzero $\kappa$ under fully symmetric conditions. Therefore, one finds an approximately symmetric mode and two complex modes whose imaginary parts determine which mode crosses the lasing threshold first.
From this model, we can also deduce that an EP3 is not generically reachable with strictly uniform $\kappa_{ij}=\kappa$ and the diagonal pattern $(i\gamma,0,-i\gamma)$. Since an EP3 requires satisfying multiple algebraic constraints, it typically demands adjusting at least two independent parameters. Therefore, to realize an EP3 in a ternary non-Hermitian system, one must carefully tune asymmetries in the couplings and/or the diagonal terms. This result is consistent with the literature, which shows that achieving EP3 requires more precise engineering than EP2.
@@ 4 Soft-wall model of EPs and the Ferrell-Glover-Tinkham sum rule
One could write the above action in a slightly different way to simulate EPs. To simulate photonic EPs holographically more accurately, one should note that the boundary CFT (or more generally, a QFT) encodes the non-Hermitian features through “complex sources” for operators, while the bulk gravity captures the emergence of EPs via coalescence in the spectrum of quasinormal modes (QNMs) or via phase transitions in the free energy, as in the case of confinement/deconfinement discussed above.
The gain and loss can be modeled holographically by the imbalances in the inflow/outflow of a conserved charge (such as a $U(1)$ global symmetry on the boundary), dual to a bulk gauge field. PT symmetry can be imposed by ensuring that the sources and their conjugates are complex mirrors (e.g., $s$ and $s^{*}$), which makes the boundary action PT-invariant. Here, we introduce a minimal model to capture the core behavior of EPs based on an AdS/QCD-inspired setup. This model is a probe backreacting scalar in $\text{AdS}_{5}$ for a $4d$ boundary theory, tunable to QCD-like scales via the bulk potential. The theory has a $U(1)$ global symmetry and a complex scalar operator $\mathcal{O}$ of dimension $\Delta=3$, which, for instance, can mimic a mesonic operator in AdS/QCD.
The action is
@@
This action is the standard dilaton-like potential in holographic QCD models, but here $\phi$ introduces non-Hermiticity.
The equations of motion that are being obtained by varying the action are
@@
Then, at the AdS boundary ($z$ $\to$ 0, where $z$ is the radial coordinate, with the UV at the boundary and the IR deep in the bulk), we can impose Dirichlet conditions for sources as $\phi(z,x^{\mu})\sim s(x^{\mu})z^{\Delta}+\langle\mathcal{O}\rangle z^{4-\Delta}+$..., and similarly for $\phi^{*}\sim s^{*}(x^{\mu})z^{\Delta}+\langle\mathcal{O}^{\dagger}\rangle z^{4-\Delta}$.
In order to have PT symmetry, we should take $s(x^{\mu})$ real and keep it even under parity $(P:x^{1}\to-x^{1})$ and time-reversal $(T:i\to-i)$, making the boundary action $\int s\mathcal{O}+s^{*}\mathcal{O}^{\dagger}$ invariant.
Non-Hermiticity enters via the “imbalance parameter”, $x=\frac{s-s^{*}}{s+s^{*}}$, which is real for PT symmetry. The gain/loss can be modeled as $|s|>0$, which drives the charge inflow/outflow. The positive $\text{Im}(s)$ acts as gain (particle creation), while negative $\text{Im}(s)$ acts as loss or absorption.
For the three-resonator system, one could make $s(x^{1})$ piecewise constant across the three spatial “sites”, (for instance $s(x^{1})=s_{0}$ for $|x^{1}|<a/3$, and varying the couplings via gradients), which mimic a lattice. However, for simplicity, one can take $s$ homogeneous and constant.
For capturing the photonic EPs, we can then study different scales. First, for $|s|<M\approx\sqrt{3/\lambda}$, which is the critical mass scale from the potential, we have the PT-unbroken phase. In this case, real bulk solutions exist, where $\phi$ and the metric are real and the null energy condition is satisfied. Also, the boundary spectrum (QNMs of $\phi$) has real frequencies $\omega_{n}$, which are dual to stable photonic modes. The Dyson map which corresponds to a complexified bulk $U(1)$ gauge shift $A_{M}\to A_{M}+i\alpha\partial_{M}\log\eta$ maps it to a unitary theory.
@@
Third, in the PT-broken phase ($|s|>M$), the complex bulk metric can be written as
@@
Note that in this case, the $4D$ strongly coupled boundary CFT is local, while the $5D$ bulk is non-local, as the gravitational degrees of freedom (metric $g_{MN}$, scalar $\phi$, and gauge field $A_{M}$) are described by a gauge theory where diffeomorphism invariance introduces non-local effects. The non-Hermiticity (gain/loss) is dual to complex bulk-boundary conditions, and the EP corresponds to a critical point in the bulk where QNMs coalesce, often tied to non-local features such as horizon instabilities or complex metric solutions that violate the NEC in the PT-broken phase. Thus, the non-locality arises from gravitational gauge redundancy and the radial smearing of interactions, with the EP manifesting as a non-local critical point in the QNM spectrum.
Then, to numerically find the EP, one needs to discretize the radial Sturm-Liouville operator and build a generalized eigenproblem. The radial equation can be written as
@@
or in operator form, $L_{0}\psi+\lambda L_{1}\psi=0$, with $L_{1}=p(z)$ multiplication. Here, $p(\lambda)=\text{det}(\mathcal{M}-\lambda I)$ is the characteristic polynomial.
Using our holographic model, we then can build a holographic block operator $H$, which is the effective non-Hermitian Hamiltonian for the ternary coupled microring system, inspired by bottom-up AdS/QCD holography, and can be written as
@@
where $L_{0}$ and $L_{1}$ are radial differential operators that come from discretizing the $\text{AdS}_{5}$ wave equation with flavor brane embedding functions. So $L_{0}$ and $L_{1}$ carry the holographic bulk physics. Each block corresponds to one ring resonator embedded as a boundary degree of freedom, dual to a bulk flavor brane. Off-diagonal terms $(\kappa,\kappa_{13})$ encode hopping/coupling between rings, which holographically correspond to mixing between bulk fields on different branes. The diagonal terms include gain and loss $(\pm i\gamma L_{1})$, which holographically represent non-Hermitian boundary conditions, meaning ingoing/outgoing flux at the AdS boundary.
Note that in the language of optics, $H$ is the effective coupled-mode Hamiltonian governing the ternary microring system, while in holographic language, $H$ is the block operator encoding the bulk-boundary dynamics of three coupled flavor branes. The real parts of its eigenvalues give the resonant frequencies, while the imaginary parts give the gain/loss rates. Exceptional points (EPs) appear where two or more eigenvalues/eigenvectors coalesce, which in the holographic picture means the bulk modes become degenerate under the non-Hermitian boundary condition. Thus, $H$ is the “bridge” between the holographic bulk description and the non-Hermitian optics experiment as via $L_{0}$ and $L_{1}$, it contains both the AdS radial structure, and via $\kappa$, the microring coupling topology.
@@ 4.1 Ferrell-Glover-Tinkham sum rule and EP in holography
The FGT sum rule states that the integrated real part of the optical conductivity over all frequencies is conserved, so
@@
In our photonic exceptional point system with a non-Hermitian Hamiltonian, the FGT sum rule can be applied analogously in the holographic dual, AdS/CFT or AdS/QCD case. The boundary optical conductivity, derived from the bulk gauge field $A_{M}$, probes the dissipative dynamics of the strongly coupled QFT dual to the photonic system. The sum rule holds across the PT-symmetry-breaking transition, ensuring that the total spectral weight of $\sigma_{1}(\omega)$ remains conserved, even as modes coalesce and eigenvalues become complex.
In the PT-unbroken phase ($|s|<M$), the conductivity has a real spectrum, akin to a normal metal. At the EP ($|s|=M$), the conductivity shows a critical redistribution of spectral weight, with modes merging. In the PT-broken phase ($|s|>M$), complex eigenvalues lead to amplified or damped modes, and the sum rule ensures that the integrated weight remains fixed, with contributions shifting to low frequencies, which is mimicking superfluid-like behavior due to gain/loss. In the holographic model, this is computed via the retarded Green’s function of the U(1) current, $G_{JJ}(\omega)\sim\sigma(\omega)$, from the bulk gauge field’s boundary response.
Figure 8: Ferrell-Glover-Tinkham sum rule for coupled ternary microcavities. This is a numerical demonstration of an FGT-style spectral-weight check for three toy pumping scenarios: even gain–gain–gain, gain–loss–loss, and anti-PT.
When one applies the FGT sum rule, the resonance positions are then set by the real parts of the eigenvalues of the holographic block operator $H$.

### v5

@@ 3.1 The holographic lasing spectra of various pumping patterns
Now, based on our model of three coupled flavor-brane gauge modes $a_{i}(z)$, we can extend the work of [26] and find analytical results for several symmetric cases which are gain-gain-gain, gain-loss-loss, and gain-gain-loss patterns, with partially pumped or evenly pumped structures, $\mathcal{PT}$-symmetric gain-loss arrangements (unbroken to broken), and then we can identify EP2/EP3.
@@
Figure 4: Two runs for finding eigenvalue trajectories in the complex plane as $\gamma$ is varied for the ternary coupled microcavity model (anti-$\mathcal{PT}$ sweep). In the left $N_{z}=20$ (grid points along holographic $z$) and in the right $N_{z}=200$.
Figure 5: Two runs for finding imaginary parts versus $\gamma$ (growth rates of modes) for the ternary coupled microcavity model (anti-$\mathcal{PT}$ sweep). In the left $N_{z}=20$ (grid points along holographic z) and in the right $N_{z}=200$.
Figure 6: Two runs for finding radial profiles of the dominant lasing eigenmode at the $\gamma$ where the growth is maximal for the ternary coupled microcavity model (anti-$\mathcal{PT}$ sweep). In the left $N_{z}=20$ (grid points along holographic z) and in the right $N_{z}=200$. One could see that similar to confining models, most of the modes are accumulated close to the wall, around $z=0$.
@@
For even pumping (gain-gain-gain), we have $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma$. For the partial pumping/gain-loss-loss, we get $\gamma_{1}=\gamma$, $\gamma_{2}=-\gamma$, $\gamma_{3}=-\gamma$, or $\gamma_{1}=\gamma$, $\gamma_{2}=\gamma$, $\gamma_{3}=0$. For the anti-$\mathcal{PT}$ ternary system, we get $\gamma_{1}=\gamma$, $\gamma_{2}=0$ and $\gamma_{3}=-\gamma$. For the $\mathcal{PT}$ pair plus a spectator case, we get a $\mathcal{PT}$ pair on modes 1 and 2, while the mode 3 is neutral or weakly coupled. So we have $\gamma_{1}=\gamma$, $\gamma_{2}=-\gamma$, $\gamma_{3}=0$.
@@
Figure 7: Phase diagram distinguishing the unbroken $\mathcal{PT}$ phase, where all eigenvalues are real, to the broken $\mathcal{PT}$ phase, where at least one eigenvalue is complex.
In figure 7, a phase diagram heatmap is shown. In the blue region (0), there is an unbroken $\mathcal{PT}$ phase where all eigenvalues are real, corresponding to symmetric lasing. In the red region (1), the $\mathcal{PT}$ symmetry is broken, with complex-conjugate eigenvalue pairs indicating symmetry breaking. This illustrates where the exceptional points occur as transitions between the phases are consistent with the physics of ternary microrings of [25].
@@ 3.2 Analytic eigenvalues for symmetric coupling
For even pumping corresponding to gain-gain-gain, where $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma$, and for the non-degenerate symmetric mode, we get $(1,1,1):\lambda_{+}=i\gamma+2\kappa$, while for the doubly degenerate orthogonal subspace with degeneracy 2, we have $\lambda_{-}=i\gamma-\kappa$.
@@
For the fully symmetric anti-$\mathcal{PT}$ diagonal pattern $(i\gamma,0,-i\gamma)$ and uniform $\kappa$, the characteristic cubic must be solved. In this case, the eigenvalues are generally distinct complex numbers, and there is no triple root for nonzero $\kappa$ under fully symmetric conditions. Therefore, one finds an approximately symmetric mode and two complex modes whose imaginary parts determine which mode crosses the lasing threshold first.
From this model, we can also deduce that an EP3 is not generically reachable with strictly uniform $\kappa_{ij}=\kappa$ and the diagonal pattern $(i\gamma,0,-i\gamma)$. Since an EP3 requires satisfying multiple algebraic constraints, it typically demands adjusting at least two independent parameters. Therefore, to realize an EP3 in a ternary non-Hermitian system, one must carefully tune asymmetries in the couplings and/or the diagonal terms. This result is consistent with the literature, which confirms that achieving EP3 requires more precise engineering than EP2.
@@ 4 Soft-wall model of EPs and the Ferrell-Glover-Tinkham sum rule
One could write the above action in a slightly different way to simulate EPs. To imitate the photonic EPs by our holographical toy model more accurately, one should note that the boundary encodes the non-Hermitian features through “complex sources” for operators, while the bulk parameters capture the emergence of EPs via coalescence in the spectrum of quasinormal modes (QNMs) or via phase transitions in the free energy, similar to confinement/deconfinement as discussed above.
The gain and loss can be modeled in this setup by the imbalances in the inflow/outflow of a conserved charge (such as a $U(1)$ global symmetry on the boundary), which is dual to a bulk gauge field. The $\mathcal{PT}$ symmetry can be imposed by ensuring that the sources and their conjugates are complex mirrors (e.g., $s$ and $s^{*}$), which makes the boundary action $\mathcal{PT}$-invariant. Here, we introduce a minimal model to capture the core behavior of EPs based on an AdS/QCD-inspired analogy. In this model, for a $4d$ boundary theory, we have a probe backreacting scalar in $\text{AdS}_{5}$, tunable to QCD-like scales via the bulk potential. The theory has a $U(1)$ global symmetry and a complex scalar operator $\mathcal{O}$ of dimension $\Delta=3$, which, for instance, can mimic a mesonic operator in AdS/QCD.
The action here is
@@
This action is the standard dilaton-like potential in holographic QCD models, but here $\phi$ introduces non-Hermiticity in the system.
The equations of motion can be obtained by varying the action as
@@
Then, at the AdS boundary ($z$ $\to$ 0, where $z$ is the radial coordinate, with the UV at the boundary and the IR deep in the bulk), we can impose Dirichlet boundary conditions for sources as $\phi(z,x^{\mu})\sim s(x^{\mu})z^{\Delta}+\langle\mathcal{O}\rangle z^{4-\Delta}+$..., and similarly for $\phi^{*}\sim s^{*}(x^{\mu})z^{\Delta}+\langle\mathcal{O}^{\dagger}\rangle z^{4-\Delta}$.
In order to have $\mathcal{PT}$ symmetry, we should take $s(x^{\mu})$ real and keep it even under parity $(P:x^{1}\to-x^{1})$ and time-reversal $(T:i\to-i)$, making the boundary action $\int s\mathcal{O}+s^{*}\mathcal{O}^{\dagger}$ invariant.
Non-Hermiticity enters via the “imbalance parameter”, $x=\frac{s-s^{*}}{s+s^{*}}$, which is real for $\mathcal{PT}$ symmetry. The gain/loss can be modeled as $|s|>0$, which drives the charge inflow/outflow. The positive $\text{Im}(s)$ acts as gain (particle creation), while negative $\text{Im}(s)$ acts as loss or absorption.
For the three-resonator system, one could make $s(x^{1})$ piecewise constant across the three spatial “sites”, (for instance $s(x^{1})=s_{0}$ for $|x^{1}|<a/3$, and varying the couplings via gradients), which mimic a lattice. However, for simplicity, one could take $s$ to be homogeneous and constant.
For capturing the photonic EPs, we can then study different scales in the system. First, for $|s|<M\approx\sqrt{3/\lambda}$, which is the critical mass scale from the potential, we have the $\mathcal{PT}$-unbroken phase. In this case, real bulk solutions exist, where $\phi$ and the metric are real and the null energy condition is satisfied. Also, the boundary spectrum (QNMs of $\phi$) has real frequencies $\omega_{n}$, which are dual to stable photonic modes. The Dyson map which corresponds to a complexified bulk $U(1)$ gauge shift $A_{M}\to A_{M}+i\alpha\partial_{M}\log\eta$ maps it to a unitary theory.
@@
Third, in the $\mathcal{PT}$-broken phase ($|s|>M$), the complex bulk metric can be written as
@@
Note that in this case, the $4D$ strongly coupled boundary CFT is local, while the $5D$ bulk is non-local, as the gravitational degrees of freedom (metric $g_{MN}$, scalar $\phi$, and gauge field $A_{M}$) are described by a gauge theory where diffeomorphism invariance introduces non-local effects. The non-Hermiticity (gain/loss) is dual to complex bulk-boundary conditions, and the EP could correspond to a critical point in the bulk where QNMs coalesce, often tied to non-local features such as horizon instabilities or complex metric solutions that violate the NEC in the $\mathcal{PT}$-broken phase. Thus, the non-locality which arises from gravitational gauge redundancy and the radial smearing of interactions with the EP, would be an analogical manifestation of a non-local critical point in the QNM spectrum.
Then, to numerically find the EP, one needs to discretize the radial Sturm-Liouville operator and build a generalized eigen-problem. The radial equation can be written as
@@
or in the operator form, $L_{0}\psi+\lambda L_{1}\psi=0$, with $L_{1}=p(z)$ multiplication. Here, $p(\lambda)=\text{det}(\mathcal{M}-\lambda I)$ is the characteristic polynomial.
Using our toy model, we then can build a “holographic-like” block operator $H$, which is the effective non-Hermitian Hamiltonian for the ternary coupled microring system, and is inspired by bottom-up AdS/QCD holography, and can be written as
@@
where $L_{0}$ and $L_{1}$ are radial differential operators that come from discretizing the $\text{AdS}_{5}$ wave equation with flavor brane embedding functions. So $L_{0}$ and $L_{1}$ carry the physics in the engineered bulk geometries. Each block corresponds to one ring resonator embedded as a boundary degree of freedom, dual to a bulk flavor brane. Off-diagonal terms $(\kappa,\kappa_{13})$ encode hopping/coupling between rings, which by the holographical analogy encode the mixing between bulk fields on different branes. The diagonal terms include gain and loss $(\pm i\gamma L_{1})$, which holographically represent non-Hermitian boundary conditions, meaning ingoing/outgoing flux at the AdS boundary.
Note that in the language of optics, $H$ is the effective coupled-mode Hamiltonian governing the ternary microring system, while in the holographic language, $H$ is the block operator which can encode the bulk-boundary dynamics of three coupled flavor branes. The real parts of its eigenvalues give the resonant frequencies, while the imaginary parts give the gain/loss rates. Exceptional points appear where two or more eigenvalues/eigenvectors coalesce, which in the holographic picture means the bulk modes become degenerate under the non-Hermitian boundary condition. Thus, $H$ is the “bridge” between the holographic bulk toy model and the non-Hermitian optical experiments, via $L_{0}$ and $L_{1}$, as it contains the AdS radial structure, and via $\kappa$, the microring coupling topology.
@@ 4.1 Ferrell-Glover-Tinkham sum rule and EP in holography
The FGT sum rule states that the integrated real part of the optical conductivity over all frequencies is conserved, so we get
@@
In our photonic exceptional point system with a non-Hermitian Hamiltonian, the FGT sum rule can be applied analogously as in the holographic dual, in AdS/CFT or in AdS/QCD toy models. The boundary optical conductivity, derived from the bulk gauge field $A_{M}$, can mimic the dissipative dynamics of the strongly coupled QFT which are dual to the photonic system. The sum rule holds across the $\mathcal{PT}$-symmetry-breaking transition, ensuring that the total spectral weight of $\sigma_{1}(\omega)$ remains conserved, even as modes coalesce and eigenvalues become complex.
In the $\mathcal{PT}$-unbroken phase ($|s|<M$), the conductivity has a real spectrum, akin to a normal metal. At the EP ($|s|=M$), the conductivity shows a critical redistribution of spectral weight, with modes merging. In the $\mathcal{PT}$-broken phase ($|s|>M$), complex eigenvalues lead to amplified or damped modes, and the sum rule ensures that the integrated weight remains fixed, with contributions shifting to low frequencies, which is mimicking superfluid-like behavior due to gain/loss. In the analogical holographic toy model, this is computed via the retarded Green’s function of the U(1) current, $G_{JJ}(\omega)\sim\sigma(\omega)$, from the bulk gauge field’s boundary response.
Figure 8: Ferrell-Glover-Tinkham sum rule for coupled ternary microcavities. This is a numerical demonstration of an FGT-style spectral-weight check for three toy pumping scenarios which are even gain-gain-gain, gain-loss-loss, and anti-$\mathcal{PT}$.
When one applies the FGT sum rule, the resonance positions are then set by the real parts of the eigenvalues of the analogical holographic block operator $H$.